An automobile manufacturer claims that their van has a 55.4 miles/gallon (MPG) rating. An independent testing firm has been contracted to test the MPG for this van. After testing 100 vans they found a mean MPG of 55.7. Assume the standard deviation is known to be 1.6. Is there sufficient evidence at the 0.01 level that the vans outperform the manufacturer's MPG rating?

Step 1 of 5: Enter the hypotheses:

Step 2 of 5: Enter the value of the z test statistic. Round your answer to two decimal places.

Step 3 of 5: Specify if the test is one-tailed or two-tailed.

Step 4 of 5: Enter the decision rule.

Step 5 of 5: Enter the conclusion.
Answer #1

Solution :

= 55.4

=55.7

S =1.6

n = 100

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :    = 55.4

Ha :     55.4

Test statistic = t

= ( - ) / S / n

= (55.7 - 55.4) / 1.6 / 100

= 1.875

Test statistic = t =  1.875

P-value = 2 * 0.0319 = 0.0636

= 0.01

P-value >

0.0636 > 0.01

Fail to reject the null hypothesis .

There is insufficient evidence to suggest that